OpenStudy (anonymous):

The point (6,-10) is on a circle with center (2,-6). Write the standard equation of the circle.

OpenStudy (anonymous):

\[(x-2)^2+(y+6)^2=r^2\] where \(r^2\) is the square of the distance between the two points

OpenStudy (anonymous):

i.e. \[r^2=(6-2)^2+(-10+6)^2\]

OpenStudy (anonymous):

which one is it?